# ChemVLM: Exploring the Power of Multimodal Large Language Models in Chemistry Area

### v4

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success and have been applied across various scientific fields, including chemistry. However, many chemical tasks require the processing of visual information, which cannot be successfully handled by existing chemical LLMs. This brings a growing need for models capable of integrating multimodal information in the chemical domain. In this paper, we introduce ChemVLM, an open-source chemical multimodal large language model specifically designed for chemical applications. ChemVLM is trained on a carefully curated bilingual multimodal dataset that enhances its ability to understand both textual and visual chemical information, including molecular structures, reactions, and chemistry examination questions. We develop three datasets for comprehensive evaluation, tailored to Chemical Optical Character Recognition (OCR), Multimodal Chemical Reasoning (MMCR), and Multimodal Molecule Understanding tasks. We benchmark ChemVLM against a range of open-source and proprietary multimodal large language models on various tasks. Experimental results demonstrate that ChemVLM achieves competitive performance across all evaluated tasks. Our model can be found at https://huggingface.co/AI4Chem/ChemVLM-26B.
Codehttps://github.com/AI4Chem/ChemVlm
Modelhttps://huggingface.co/AI4Chem/ChemVLM-26B
@@ Introduction
To address this issue, we introduce ChemVLM, an open-source multimodal large language model specifically designed for the chemical domain, built upon the ViT-MLP-LLM architecture (Liu et al. 2023). Our model combines a state-of-the-art Vision Transformer (ViT) (Dosovitskiy et al. 2020) for robust image encoding with ChemLLM (Zhang et al. 2024a), a domain-specific Large Language Model trained on billions of tokens from high-quality chemical data. This architecture effectively bridges the gap in domain-specific knowledge needed to process multimodal chemical tasks. A Multi-Layer Perceptron (MLP) layer is adopted as a projector to perform image-text alignment. We implement a two-stage supervised fine-tuning strategy to enhance model performance. In the first stage, we freeze the LLM component to focus on modality alignment. In the second stage, we unfreeze and update all parameters to adapt the model to downstream tasks. This structured approach ensures that ChemVLM achieves competitive performance across a variety of chemical tasks.
@@
Figure 1: Overall architecture and two-stage training of ChemVLM. ChemVLM combines the advantage of an advanced vision transformer and a large language enriched with chemical knowledge, ensuring the strong ability of multimodal chemical knowledge understanding and reasoning.
@@
Overview of Multimodal Large Language Models. In addition to textual content, visual information encompasses a wealth of data, and integrating these modalities through large language models (MLLMs) has demonstrated superior comprehension, reasoning, and generation capabilities compared to traditional models. Notable MLLMs include GPT-4V (OpenAI 2023), which extends GPT-4 with a visual module, and the Gemini series (Team et al. 2023; Reid et al. 2024), which processes audio, video, and text inputs with enhanced long-context understanding. Proprietary models like QwenVL-Plus/Max (Bai et al. 2023a), MM 1 (McKinzie et al. 2024), and Claude-3V (Anthropic 2024) also exhibit exceptional multimodal performance. Open-source MLLMs, such as the LLaVA series (Liu et al. 2023, 2024), VisionLLM (Wang et al. 2024), CogVLM (Wang et al. 2023b), LLaMA-Adapter V2 (Gao et al. 2023), and ShareGPT4V (Chen et al. 2023), continue to evolve but still lag behind their proprietary counterparts in performance.
Overview of Chemical OCR and MMCR. Various ways have been attempted by prior works to extract visual information from the chemical images. Works on diagram parsing focused on the segmentation of molecular images (Joshua et al. 2019; EJ Beard 2020) and the recognition of their chemical structures (Qian et al. 2023b; IV Filippov 2009; K Rajan 2020). Some researchers want to understand the relationships between the molecules, i.e., reaction schemes. Wilary and Cole proposed ReactionDataExtractor (DM Wilary 2021) to extract reaction schemes from the diagrams. A method called Chemgrapher (M et al. 2020) suggests dealing with the problem in a modular fashion by using a segmentation algorithm to segment the images containing chemical graphs to detect atom locations, bonds, and charges. Schrödinger group presented an image-to-SMILES generation model (J et al. 2019): a convolutional neural network encodes the input molecular image into a hidden representation, and a recurrent neural network decodes the SMILES string of the molecule. Only a few try to predict the properties of molecules. Chithrananda et al. make one of the first attempts to systematically evaluate transformers on Multimodal Molecule Understanding tasks such as molecular property prediction (Seyone Chithrananda 2020). We provide an MLLM method to help solve the problem of chemical OCR and can work as a chemical properties answer assistant.
@@
As shown in Figure 1, ChemVLM follows the architecture from LLaVA (Liu et al. 2023, 2024) in the fashion of "ViT-MLP-LLM". This framework integrates a Vision Transformer (ViT) followed by Multi-Layer Perceptron (MLP) components, seamlessly connected to a Large Language Model (LLM). We adopt InternViT-6B (Chen et al. 2024c) as the vision encoder and ChemLLM-20B (Zhang et al. 2024a) as the language model, as it demonstrates strong chemical capabilities. The projector is an MLP to convert the visual feature into the language embedding space, whose weights are randomly initialized. The visual encoder takes images of $448\times 448$ resolution as input and extracts high-level features from them. The visual features are then transformed into the same dimension of the language tokens, aligning them with the text space. Simultaneously, the textual input passes through the tokenizer to be converted into a sequence of tokens. The image and text tokens are then concatenated to form a unified token sequence and then fed into the LLM. The LLM synthesizes the multimodal tokens to generate a cohesive and contextually relevant response. This pipeline enables the model to perform comprehensive analyses of the input, thereby enhancing capability in complex chemical environments.
@@
This section details our approach to enhancing our model’s recognition and understanding of images of molecules and reactions, and multimodal reasoning tasks in chemistry. We aim to improve the model’s capability in real-world multimodal scenarios by leveraging diverse datasets and educational resources. We visualize the data preparation work in Figure 2.
@@
  • Hand-drawn style. We utilize the DECIMER HDM (Brinkhaus et al. 2023) dataset, which containing over 7,000 hand-drawn molecular structure images. This dataset includes drawings created with different inks and types of pens.
@@
  • Electronic document style. This subset is characterized by images generated by chemical drawing tools, including ChemDraw (Li et al. 2004), RDKit (Landrum 2014) and Indigo***A universal cheminformatics toolkit, utilities and database search tools can be found in https://github.com/epam/Indigo. The images are created with varying line styles and color presets.
@@
Existing multimodal chemical models (Qian et al. 2023b; IV Filippov 2009; K Rajan 2020; Qian et al. 2023a; Zhang et al. 2024b) primarily focused on modality conversion tasks, such as OCR and molecule captioning, and have achieved a remarkable progress. However, with large language models increasingly being used as tools for knowledge reasoning and scientific discovery (Mirza et al. 2024), the need for a deeper, more integrated understanding of multimodal knowledge has become critical. Some multimodal question-answering datasets (Mirza et al. 2024; Luo et al. 2023) have been involved in the field of chemistry, however, these efforts often focus solely on evaluating the model’s ability to recognize and statistically analyze local structures or sub-patterns within non-text modalities, such as identifying specific functional groups within molecular structures. This approach typically results in weaker logical reasoning capabilities, which are insufficient for real-world multimodal chemistry reasoning tasks. For tasks like solving chemistry problems in high school chemistry exams, the model needs to possess strong logical reasoning skills, be able to capture critical information from the questions, and combine it with extensive domain-specific knowledge to provide accurate and reasonable answers.
Given the importance of applying chemical knowledge comprehensively and the understanding of real-world multimodal data, we utilize examination data within the chemistry discipline to build a Q&A dataset. Inspired by C-MHChem (Zhang et al. 2024a), we carefully curate and deduplicate a dataset containing 200,000 high-quality multimodal chemistry questions from the OpenDataLab platform (He et al. 2024a), covering Chinese education from secondary school to graduate levels. It includes various problem types such as multiple choice, fill-in-the-blank, and short answer. The questions are designed to assess diverse skills of test-takers, including error correction, knowledge-based Q&A, complex reasoning, and experiment protocol design. This dataset aims to equip the model with a deeper understanding of chemistry, enabling it to tackle complex multimodal challenges in real-world scenarios.
@@
Model&Method Params Avg Sim. (%↑) Tani@1.0 (%↑)
Decimer (Rajan, Zielesny, and Steinbeck 2021) EfficientNet-B7 (66M) + transformer 85.0 77.3
Molscribe (Qian et al. 2023b) Swin-B (88M) + 6-layer transformer 92.0 89.1
Qwen-VL-Chat (Bai et al. 2023a) 9.6B 5.0 0.0
LLaVA-v1.5-13B (Liu et al. 2023) 13B 1.0 0.0
InternVL-v1.5 (Chen et al. 2024c) 26B 9.0 0.0
Yi-VL-Plus (Young et al. 2024) 34B 5.0 0.0
GPT-4V (OpenAI 2023) 15.0 2.1
ChemVLM (ours) 26B 71.0 42.9
Table 1: Results on ChemOCR. We compare our ChemVLM with models specially designed for chemical OCR task: Decimer and Molscribe, and other MLLMs: Qwen-VL-Chat, Yi-VL-Plus, InternVL-v1.5, and GPT-4V. Tanimoto similarities are written as Avg Sim. (↑), and tanimoto@1.0 written as Tani@1.0 (%↑). High is better for both metrics.
@@
In this section, we conduct extensive experiments to assess the chemical capabilities of ChemVLM and other competing models on benchmarks of various tasks. Here we show the overall performance in Figure 3.
Figure 3: Overall performance of models.
Model&Method Params ScienceQA Chemistry(%↑) CMMU Chemistry(%↑) MMCR-Bench (%↑)
Qwen-VL-Chat (Bai et al. 2023a) 9.6B 65.3 22.1 28.3
LLaVA-v1.5-13B (Liu et al. 2023) 13B 43.6 18.6 23.1
InternVL-v1.5 (Chen et al. 2024c) 26B 74.0 26.9 36.5
Yi-VL-Plus (Young et al. 2024) 34B 68.1 29.1 35.3
GPT-4V (OpenAI 2023) 71.9 24.2 40.1
ChemVLM (ours) 26B 71.2 31.6 41.7
@@
To assess the multimodal capabilities of models on chemical tasks, we utilize benchmark datasets both the publicly available and our contributed. Our evaluation focuses on the three critical chemical tasks, including Chemical Optical Character Recognition (Chemical OCR), Multimodal Chemical Reasoning (MMCR), Multimodal Molecule Understanding.
We utilize the chemistry part of open-source multimodal benchmarks, CMMU (He et al. 2024b) and ScienceQA (Lu et al. 2022) for MMCR evaluation. To evaluate model’s capability of Multimodal Molecule Understanding, we construct a multimodal chemical benchmark named MMChemBench, derived from ChemBench (Zhang et al. 2024a), including two image-text tasks: molecule caption and molecular property prediction.
Moreover, we build two datasets for a more comprehensive benchmarking called ChemOCR and MMCR-Bench. Specifically, ChemOCR includes 1000 chemical OCR image-text pairs collected from an open-source chemical database. And MMCR-Bench includes 1000 high-quality multimodal chemistry exam questions from the Chinese college entrance examination.
@@
Model&Method Params Score on molecule caption(%) Score on property prediction(%)
ChemLLM* (Zhang et al. 2024a) 20B 92.6 72.2
GPT-4* (OpenAI 2024a) 96.3 68.4
Qwen-VL-Chat (Bai et al. 2023a) 9.6B 86.6 56.1
LLaVA-v1.5-13B (Liu et al. 2023) 13B 75.9 19.6
InternVL-v1.5 (Chen et al. 2024c) 26B 78.7 32.2
Yi-VL-Plus (Young et al. 2024) 34B 89.3 27.4
GPT-4V (OpenAI 2023) 95.2 38.6
ChemVLM (ours) 26B 98.2 80.9
Table 3: Results on MMChemBench. We report the total score using the method of calculating a score for single-choice problems. Notably, we report ChemLLM-20B’s and GPT-4’s original performance on MMChemBench’s corresponding data source: ChemBench, with text-only input. ’*’ means this is a text-only LLM.
(a) Chinese version
(b) English version
Figure 4: A qualitative comparison of answers on MMCR-Bench on GPT-4V and our ChemVLM. Mistakes within the answers are highlighted in red, whereas detailed and accurate parts are emphasized in green. Since this is a Chinese exam question, we prepare the original Chinese text and the English translation of it. This shows the strong MMCR capability of ChemVLM.
@@
An answer gets one point when it matches all the right choices of a single or multiple-choice problem, else it gets zero point.
@@
Results on the MMCR tasks are summarized in Table 2. Our model achieves state-of-the-art (SOTA) for the chemical part on the CMMU benchmark. Additionally, ChemVLM demonstrates strong performance in the chemistry section of ScienceQA, closely matching that of GPT-4V. Upon analyzing the ScienceQA dataset, we find that many of the questions rely on common sense, making them easier for MLLMs to solve. When it comes to real exam questions, ChemVLM shows more promising performance, which surpasses GPT-4V by 1.7% and achieves state-of-the-art performance.
@@
We present MMChemBench, an extension of ChemBench, designed to evaluate our models’ performance in Multimodal Molecule Understanding tasks which include molecule caption and molecule property prediction with multimodal information input. We also compare ChemVLM with ChemLLM to assess the impact of incorporating visual modality into the model. The results are presented in Table 3, which demonstrate that ChemVLM achieves state-of-the-art performance in both tasks. A significant improvement is observed in the property prediction task compared to other MLLMs. The findings highlight that incorporating visual information greatly enhances the model’s ability to understand domain-specific knowledge, such as molecular structures.
Figure 5: A qualitative comparison of answers on ChemOCR on GPT-4V and our ChemVLM. Mistakes within the answers are highlighted in red, whereas detailed and accurate parts are emphasized in green.
@@
Quantitative results from the previous sections have showcased ChemVLM’s superior chemical capabilities. In this part, we present qualitative results to compare the outputs of our model with those from a strong proprietary MLLM, GPT-4V, as shown in Figure 4 and Figure 5. The results indicate that our model exhibits a better understanding of the questions and provides more accurate answers for chemical OCR and examination questiones. This highlights ChemVLM’s proficiency in leveraging chemical knowledge to address complex problems.
@@
In this study, we introduce ChemVLM, an open-source, multimodal model tailored for applications in chemistry, leveraging the strengths of ChemLLM-20B and InternViT-6B through a two-stage training process. Our evaluations indicate that ChemVLM exceeds most existing methods, significantly impacting the AI and chemistry fields.
Currently, ChemVLM primarily incorporates image modalities and lacks processing capabilities for molecular graph and time-series data, limiting its effectiveness in capturing complex chemical phenomena. Additionally, the high computational costs associated with its training and deployment pose barriers to widespread adoption.
Looking forward, we aim to enhance ChemVLM by adding new modalities, such as graph and time-series data, and by developing a range of models with varying parameter sizes to suit diverse scenarios. We also plan to explore more efficient training methods and develop compact models to reduce costs and facilitate broader deployment, thus advancing LLM’s role in chemistry research.

### v5

@@ Abstract
Large Language Models (LLMs) have achieved remarkable success and have been applied across various scientific fields, including chemistry. However, many chemical tasks require the processing of visual information, which cannot be successfully handled by existing chemical LLMs. This brings a growing need for models capable of integrating multimodal information in the chemical domain. In this paper, we introduce ChemVLM, an open-source chemical multimodal large language model specifically designed for chemical applications. ChemVLM is trained on a carefully curated bilingual multimodal dataset that enhances its ability to understand both textual and visual chemical information, including molecular structures, reactions, and chemistry examination questions. We develop three datasets for comprehensive evaluation, tailored to Chemical Optical Character Recognition (OCR), Multimodal Chemical Reasoning (MMCR), and Multimodal Molecule Understanding tasks. We benchmark ChemVLM against a range of open-source and proprietary multimodal large language models on various tasks. Experimental results demonstrate that ChemVLM achieves competitive performance across all evaluated tasks.
Codehttps://github.com/lijunxian111/ChemVlm
@@ Introduction
To address this issue, we introduce ChemVLM***This work was done during their internship at Shanghai Artificial Intelligence Laboratory. , an open-source multimodal large language model specifically designed for the chemical domain, built upon the ViT-MLP-LLM architecture (Liu et al. 2023). Our model combines a state-of-the-art Vision Transformer (ViT) (Dosovitskiy et al. 2020) for robust image encoding with ChemLLM (Zhang et al. 2024a), a domain-specific Large Language Model trained on billions of tokens from high-quality chemical data. This architecture effectively bridges the gap in domain-specific knowledge needed to process multimodal chemical tasks. A Multi-Layer Perceptron (MLP) layer is adopted as a projector to perform image-text alignment. We implement a two-stage supervised fine-tuning strategy to enhance model performance. In the first stage, we freeze the LLM component to focus on modality alignment. In the second stage, we unfreeze and update all parameters to adapt the model to downstream tasks. This structured approach ensures that ChemVLM achieves competitive performance across a variety of chemical tasks.
@@
Figure 1: Overall architecture of ChemVLM. ChemVLM combines the advantage of an advanced vision transformer and a large language enriched with chemical knowledge, ensuring the strong ability of multimodal chemical knowledge understanding and reasoning.
@@
Overview of Multimodal Large Language Models. In addition to textual content, visual information encompasses a wealth of data, and integrating these modalities through large language models (MLLMs) has demonstrated superior comprehension, reasoning, and generation capabilities compared to traditional models. Notable MLLMs include GPT-4V (OpenAI 2023), which extends GPT-4 with a visual module, and the Gemini series (Team et al. 2023; Reid et al. 2024), which processes audio, video, and text inputs with enhanced long-context understanding. Proprietary models like QwenVL-Plus/Max (Bai et al. 2023a), MM1 (McKinzie et al. 2024), and Claude-3V (Anthropic 2024) also exhibit exceptional multimodal performance. Open-source MLLMs, such as the LLaVA series (Liu et al. 2023, 2024), VisionLLM (Wang et al. 2024a), CogVLM (Wang et al. 2023b), LLaMA-Adapter V2 (Gao et al. 2023), and ShareGPT4V (Chen et al. 2023), continue to evolve but still lag behind their proprietary counterparts in performance.
Overview of Chemical OCR and MMCR. Various ways have been attempted by prior works to extract visual information from the chemical images. Works on diagram parsing focused on the segmentation of molecular images (Joshua et al. 2019; EJ Beard 2020) and the recognition of their chemical structures (Qian et al. 2023b; IV Filippov 2009; K Rajan 2020). Some researchers want to understand the relationships between the molecules, i.e., reaction schemes. Wilary and Cole proposed ReactionDataExtractor (DM Wilary 2021) to extract reaction schemes from the diagrams. A method called Chemgrapher (M et al. 2020) suggests dealing with the problem in a modular fashion by using a segmentation algorithm to segment the images containing chemical graphs to detect atom locations, bonds, and charges. Only a few try to predict the properties of molecules. Chithrananda et al. make one of the first attempts to systematically evaluate transformers on Multimodal Molecule Understanding tasks such as molecular property prediction (Seyone Chithrananda 2020). We provide an MLLM method to help solve the problem of chemical OCR and can work as a chemical properties answer assistant.
@@
As shown in Figure 1, ChemVLM follows the architecture from LLaVA (Liu et al. 2023, 2024) in the fashion of "ViT-MLP-LLM". This framework integrates a Vision Transformer (ViT) followed by Multi-Layer Perceptron (MLP) components, seamlessly connected to a Large Language Model (LLM). We adopt InternViT-6B (Chen et al. 2024c) as the vision encoder and ChemLLM-20B (Zhang et al. 2024a) as the language model, as it demonstrates strong chemical capabilities. The projector is an MLP to convert the visual feature into the language embedding space, whose weights are randomly initialized. The visual encoder takes images of $448\times 448$ resolution as input and extracts high-level features from them. The visual features are then transformed into the same dimension of the language tokens, aligning them with the text space. Simultaneously, the textual input passes through the tokenizer to be converted into a sequence of tokens. The image and text tokens are then concatenated to form a unified token sequence and then fed into the LLM. The LLM synthesizes the multimodal tokens to generate a cohesive and contextually relevant response. This pipeline enhances the model’s capability in complex chemical environments.
@@
This section details our approach to enhancing our model’s recognition and understanding of images of molecules and reactions, and multimodal reasoning tasks in chemistry. We aim to improve the model’s capability in real-world multimodal scenarios by leveraging diverse datasets and educational resources. We visualize the data preparation work in Figure 3. The data production process is as follows. (1). We collect raw data encompassing various tasks from reliable open-source chemical datasets and databases. (2). We apply diverse image transformation techniques for data augmentation, including style transformations (e.g., handwritten styles, graph paper backgrounds), rotations, blurring, and other enhancement techniques. (3). A variety of templates are carefully designed to cover different task scenarios and improve diversity (please refer to the Appendix for template details). (4). These templates are filled with raw data to create high-quality QA pairs, which will be used to train and evaluate MLLMs. Here we show the data distribution of our train data and benchmarks in Figure 2. Detailed explanation can be found in the Appendix.
Figure 2: Data distribution of our train data and benchmarks.
@@
  • Hand-drawn style. We utilize the DECIMER HDM (Brinkhaus et al. 2023) dataset that contains over 7,000 hand-drawn molecular structure images. This dataset includes drawings created with different inks and types of pens.
@@
  • Electronic document style. This subset is characterized by images generated by chemical drawing tools, including ChemDraw (Li et al. 2004), RDKit (Landrum 2014) and IndigoA universal cheminformatics toolkit, utilities and database search tools can be found in https://github.com/epam/Indigo. The images are created with varying line styles and color presets.
@@
Existing multimodal chemical models (Qian et al. 2023b; IV Filippov 2009; K Rajan 2020; Qian et al. 2023a; Zhang et al. 2024b) primarily focused on modality conversion tasks, such as OCR and molecule captioning, and have achieved a remarkable progress. However, with large language models increasingly being used as tools for knowledge reasoning and scientific discovery (Mirza et al. 2024), the need for a deeper, more integrated understanding of multimodal knowledge has become critical. Some multimodal question-answering datasets (Mirza et al. 2024; Luo et al. 2023) have been involved in the field of chemistry, however, these efforts often focus solely on evaluating the model’s ability to recognize and statistically analyze local structures or sub-patterns within non-text modalities, such as identifying specific functional groups within molecular structures. For tasks like solving chemistry problems in high school chemistry exams, the model needs to possess strong logical reasoning skills, be able to capture critical information from the questions, and combine it with extensive domain-specific knowledge to provide accurate and reasonable answers.
Given the importance of the problem above, we utilize examination data within the chemistry discipline to build a Q&A dataset. Inspired by C-MHChem (Zhang et al. 2024a), we carefully curate and deduplicate a dataset containing 200,000 high-quality multimodal chemistry questions from the OpenDataLab platform (He et al. 2024a), covering Chinese education from secondary school to graduate levels§§§The chemistry QA pairs from OpenDataLab are originally in Chinese, and we translate 10,704 of them into English using the Bing translator from pypi/translators.. It includes various problem types such as multiple choice, fill-in-the-blank, and short answer. The questions are designed to assess diverse skills of test-takers, including error correction, knowledge-based Q&A, complex reasoning, and experiment protocol design. This dataset aims to equip the model with a deeper understanding of chemistry, enabling it to tackle complex multimodal challenges in real-world scenarios.
@@
Model&Method Avg Sim. (%↑) Tani@1.0 (%↑)
Decimer 85.0 77.3
Molscribe 92.0 89.1
Qwen-VL-Chat 5.0 0.0
LLaVA-v1.5-13B 1.0 0.0
InternVL-v1.5 9.0 0.0
Yi-VL-Plus 5.0 0.0
GPT-4V 15.0 2.1
ChemVLM-26B (ours) 71.0 42.9
Table 1: Results on ChemOCR. Tanimoto similarities are written as Avg Sim. (%↑), and Tanimoto@1.0 written as Tani@1.0 (%↑). High is better for both metrics.
@@
In this section, we conduct extensive experiments to assess the chemical capabilities of ChemVLM and other competing models on benchmarks of various tasks.
Model ScienceQA CMMU MMCR-
&Method Chem (%↑) Chem (%↑) Bench (%↑)
Qwen-VL-Chat 65.3 22.1 28.3
LLaVA-v1.5-13B 43.6 18.6 23.1
InternVL-v1.5 74.0 26.9 36.5
Yi-VL-Plus 68.1 29.1 35.3
GPT-4V 71.9 24.2 40.1
ChemVLM-26B (ours) 71.2 31.6 41.7
@@
To assess the multimodal capabilities of models on chemical tasks, we evaluate their performance on both publicly available and self-curated benchmarks. Our evaluation focuses on the three critical chemical tasks, including Chemical Optical Character Recognition (Chemical OCR), Multimodal Chemical Reasoning (MMCR), and Multimodal Molecule Understanding.
We utilize the chemistry part of open-source multimodal benchmarks, CMMU (He et al. 2024b) and ScienceQA (Lu et al. 2022) for MMCR evaluation. To evaluate the model’s capability of Multimodal Molecule Understanding, we construct a multimodal chemical benchmark named MMChemBench, derived from ChemBench (Zhang et al. 2024a) with two image-text tasks: molecule caption and molecular property prediction. Meanwhile, we employ other scientific disciplines within the CMMU (He et al. 2024b) framework to assess ChemVLM’s generalization capabilities in different areas. We also adopt Scibench (Wang et al. 2024b), a college-level chemistry test set consisting of purely textual data, to evaluate the model’s problem-solving abilities in more complex tasks.
Moreover, we build two datasets for a more comprehensive benchmarking called ChemOCR and MMCR-Bench. Specifically, ChemOCR includes 1000 chemical OCR image-text pairs collected from an open-source chemical database (Qian et al. 2023b). And MMCR-Bench includes 1000 high-quality multimodal chemistry exam questions from the Chinese college entrance examination.
@@
Figure 4: In the left figure, we compare ChemVLM with three other MLLMs on other subjects aside from chemistry on CMMU. In the right figure, we show results on the subsets related to chemistry on Scibench. The numbers represent the performance of ChemVLM.
Figure 5: A qualitative comparison of answers on MMCR-Bench between GPT-4V and our ChemVLM. Mistakes within the answers are highlighted in red, whereas detailed and accurate parts are emphasized in green. Since this is a Chinese exam question, we prepare the original Chinese text and the English translation of it. This shows the strong MMCR capability of ChemVLM.
@@
An answer gets one point when it matches all the right choices of a multiple-choice problem, else it gets zero point.
@@
Results on the MMCR tasks are summarized in Table 2. Our model achieves state-of-the-art (SOTA) for the chemical part on the CMMU benchmark. Additionally, ChemVLM demonstrates strong performance in the chemistry section of ScienceQA, closely matching that of GPT-4V. Upon analyzing the ScienceQA dataset, we find that many of the questions rely on common sense, making them easier for MLLMs to solve. When it comes to MMCR-bench, ChemVLM shows more promising performance, which surpasses GPT-4V by 1.7% and achieves state-of-the-art performance.
Model molecule property
&Method caption(%) prediction(%)
ChemLLM* 92.6 72.2
GPT-4* 96.3 68.4
Qwen-VL-Chat 86.6 56.1
LLaVA-v1.5-13B 75.9 19.6
InternVL-v1.5 78.7 32.2
Yi-VL-Plus 89.3 27.4
GPT-4V 95.2 38.6
ChemVLM-26B (ours) 98.2 80.9
Table 3: Results on MMChemBench. We report the total score using the method of calculating a score for choice problems. Notably, we report text-only LLMs’(*) performance on MMChemBench’s corresponding data source: ChemBench.
@@
We present MMChemBench, an extension of ChemBench, designed to evaluate our models’ performance in Multimodal Molecule Understanding tasks which include molecule caption and molecule property prediction. We also compare ChemVLM with ChemLLM and GPT-4 to assess the impact of incorporating visual modality into the model. The results are presented in Table 3, which demonstrate that ChemVLM achieves state-of-the-art performance in both tasks. A significant improvement is observed in the property prediction task compared to other MLLMs. The findings highlight that incorporating visual information greatly enhances the model’s ability to understand domain-specific knowledge, such as molecular structures.
Figure 6: A qualitative comparison of answers on ChemOCR on GPT-4V and our ChemVLM. Red and green parts are the same as Figure 5.
Study on Multidisciplinary Competence
Our proposed ChemVLM is primarily trained on a proprietary chemistry dataset. To evaluate the model’s generalization capabilities, we tested it on test sets from other scientific fields within the CMMU framework. As shown in Figure 4, the experimental results indicate that ChemVLM performs exceptionally well in other disciplines, surpassing renowned multimodal LLMs and InternVL-v1.5-26B. This suggests that large-scale training on high-quality problems enables our model to exhibit remarkable capabilities across different scientific domains.
Study on More Difficult Questions
To test the model’s reasoning capability on complex chemistry problems, we evaluated various models using the Scibench (Wang et al. 2024b) dataset. Scibench is a text-only benchmark designed for university-level science questions extracted from exercise books, including topics like quantum chemistry and physical chemistry. As shown in Figure 4, our model outperformed all models mentioned in their paper, except for the GPT-4 series. Notably, in the "Chemmic" and "Matter" subtasks, our model achieved better results than GPT-4. This highlights ChemVLM’s reasoning ability on purely textual, more difficult questions, despite being designed to address multimodal chemistry problems.
@@
In this part, we present qualitative results to compare the outputs of our model with those from a strong proprietary MLLM, GPT-4V, as shown in Figure 5 and Figure 6. The results indicate that our model exhibits a better understanding of the questions and provides more accurate answers for chemical OCR and examination questions. This highlights ChemVLM’s proficiency in leveraging chemical knowledge to address multimodal problems.
@@
In this study, we introduce ChemVLM, an open-source, multimodal model tailored for applications in chemistry. Comprehensive evaluations indicate that ChemVLM exceeds most existing methods, advancing the AI and chemistry fields.
Currently, ChemVLM primarily incorporates image modalities and lacks processing capabilities for molecular graph and time-series data, limiting its effectiveness in capturing complex chemical phenomena. Additionally, the high computational costs may hinder its wider adoption.
In the future, we aim to enhance ChemVLM by adding new modalities, such as graph and time-series data. We also plan to explore more efficient training methods and develop a range of models with varying parameter sizes to suit diverse scenarios, thereby advancing LLM’s role in chemistry research.
Acknowledgements
This work is supported by Shanghai Artificial Intelligence Laboratory.